PARTRIDGE-SCHWENKE ROVIBRATIONAL ENERGY LEVELS FOR WATER: SPECTROSCOPIC AS-SIGNMENT AND COMPARISON WITH EXPERIMENTAL ENERGY LEVELS

\underline{SERGEI A. TASHKUN}, TATIANA A. PUTILOVA and SEMEN MIKHAILENKO, Laboratory of Theoretical Spectroscopy, Institute of Atmospheric Optics, 1, Akademicheskii Av., 634055 Tomsk, Russia.

Unambiguous rovibrational spectroscopic assignment (v_{1}, v_{2}, v_{3}, J, K_{a}, K_{c}) up to J,K_{a}=30 of PS energy levels for the (000), (010), (100), (020) and (001) vibrational states of nine isotopic species (H_{2}^{16}O, H_{2}^{18}O, H_{2}^{17}O, HD^{16}O, HD^{18}O, HD^{17}O, D_{2}^{16}O, D_{2}^{18}O, D_{2}^{17}O) of the water molecule will be presented. PS levels were calculated with the help of an accurate isotope dependent potential energy surface by Partridge and Schwenke^{a} and a large basis set. PS levels will be compared with experimental levels collected from the literature. Differences between PS and BT2^{b} calculated levels for the H_{2}^{16}O isotopic species will be also given and discussed.

aH. Partridge and D.W. Schwenke, J. Chem. Phys. \underline{106}(11), 4618-4639 (1997)

bR.J. Barber J. Tennyson, G.J. Harris and R.N. Tolchenov, Mon. Not. R. Astron. \underline{368}, 1087-1094 (2006)